OpenStudy (anonymous):

Given the vectors A (m, 1, 0) , B(m-1, 2m, 2) e C(1,3,-1). They are vertex of a rectangle triangle. find the value for m variable in the way that the triangle ABC be rectangle in the A angle. And so find the area of this triangle

OpenStudy (amistre64):

rectangle triangle sounds like a abomination of science

OpenStudy (amistre64):

do you mean right triangle?

OpenStudy (amistre64):

<m,1,0> <1,3,-1> --------- m+3-1 = 0 m = -2 ?

OpenStudy (amistre64):

the distances magnitudes of the vectors are the lengths of the sidses of the triangle

OpenStudy (amistre64):

1+1+9 = 11 C = sqrt(11) 5,1,0 = 6 A = sqrt(6)

OpenStudy (amistre64):

B(m-1, 2m, 2) = (-3,-4,2) = 9+16+4 = 29 B = sqrt(29) i wonder if these make a right triangle :)

OpenStudy (amistre64):

class starting soon, gotta get :)